# Classical periodic orbits in extended phase space for spherical harmonic oscillator with spin-orbit coupling

### v1

@@ IV.3 Bridge between circular orbits
The periodicity condition is now considered. There are three independent angular frequencies; $\omega_{p}$ for the spin precession, and $\omega\pm\kappa l$ for the orbital motion in $z$ direction. The frequencies for the motion in $x$ and $y$ directions ($\omega\pm\kappa l\pm\omega_{p}$) are given by the combinations of the above three. Let us first consider the case where one of the coefficients $Z_{\pm}$ vanishes. In such case, only one of the two frequencies $\omega\pm\kappa l$ should be considered in the periodicity condition. Taking into account the relation (62), let us put
$Z_{-}=0,\quad Z_{+}=\frac{s_{\perp}}{\sqrt{M\omega l}}e^{i\gamma}$ (63)
@@
$\omega+\kappa l=\frac{n}{m}\omega_{p}$ (64)
@@
Figure 4: Plot of the energy $E$ (in units of $\hbar\omega$) vs. period $T$ (in units of $2\pi/\omega$) of the periodic orbits. The vertical lines represent the diametric orbits (dotted line) and two types of circular orbits $C_{\pm}$ (dashed lines). Open circles indicate the bifurcation points (32), and the curves connected them represent the bridge orbits $(n,m)$.
@@
Each bridge orbit forms a 3-parameter family since the orientation of the total angular momentum ${\bm{j}}$ and the phase $\gamma$ in Eq. (63) are arbitrary. Figure 6 displays some examples of bridge orbits $(n,m)$ for several values of $s_{z}$. As the spin vector tilts, the orbital ripples develop and eventually subside toward spin flip.
Figure 6: Some bridge orbits $(n,m)$ with $n-m=1$, calculated for several values of $s_{z}$ (in units of $\hbar$) between $s$ and $-s$. The parameters $\kappa=0.1\,\omega/\hbar$ and $s=\hbar/2$ are used. Note that the vertical magnification is set approximately 10 times the horizontal one.
@@ IV.4 Secondary bifurcation
Each bridge orbit $(n,m)$ on the way from $C_{+}$ to $C_{-}$, the period $2n\pi/(\omega+\kappa l)$ never misses matching the middle of those for $C_{\pm}$, namely, a multiple of $2\pi/\omega$. Since $\omega+\kappa l$ becomes commensurable with $\omega$ at that point, it is also commensurable with $\omega-\kappa l$ and the combination of the modes $\mu=\omega\pm\kappa l$
@@
The orbit bifurcated from the bridge $(2,1)$ are displayed in Fig. 7 for several values of energy.
@@
$E_{\rm bif}^{(n,m,k)}=\frac{n-k}{k}\frac{\omega^{2}}{\kappa}+\frac{1}{2}\kappa s^{2}.$ (74)
and the orbit exists for $E\geq E_{\rm bif}$. Each orbit $(n,m,k)$ forms a 4-parameter family since the orientation of ${\bm{j}}$ and the arguments of the complex coefficients $Z_{\pm}$ are arbitrary, and As displayed in the $E$-$T$ plots in Fig. 8, POs with the same period $T=2\pi k/\omega$ but having different geometries $(n,m,k)$ successively emerge with increasing energy.
@@
V Summary
Classical motion was investigated in the extended phase space for the system of a spherical harmonic oscillator with simple spin-orbit coupling. General periodic solutions to the equations of motion were obtained and the bifurcation scenarios were fully clarified in both generic and specific values of the spin-orbit parameter. For arbitrary value of the spin-orbit parameter, there are two-parameter families of diametric and circular POs with frozen spin. There are two types of circular families C± in which spin is parallel or antiparallel to the orbital angular momentum. These circular orbits encounter successive bifurcations with increasing energy and generate three-parameter bridge PO families. On its way from one circular PO to another, each bridge encounters another bifurcation and the four-parameter PO family emerges, which has a constant period and survives for any higher energy. For specific values of spin-orbit parameter, different types of periodic solutions are also possible, for both frozen and moving spins.
The Berry-Tabor formula for the system under consideration suggests 4-parameter family of POs with period $T_{\pm}$, but the 4-parameter families found in this work has period $T_{0}$. The enhanced contribution of $T_{\pm}$ orbits might be explained as the effect of successive bifurcations of the circular POs and may include the contribution of bridge orbits whose periods are between $T_{+}$ and $T_{-}$. Concerning to the reduced contribution of the 4-parameter family with period $T_{0}$, cancellation may occur among the orbits proliferating with increasing energy and having different geometries. The trace formula in extended phase space [38] combined with the uniform approximations may solve these problems, that is left for a future subject.
Appendix A Berry-Tabor Formula
From the eigenvalue spectrum (5), densities of states $g^{(\pm)}$ for $\mathring{\jmath}=\mathring{l}\pm 1/2$ levels are given by
$g^{(+)}(E)$ $=\sum_{n=0}^{\infty}\sum_{\mathring{l}=0}^{\infty}(2\mathring{l}+2)\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}+\tfrac{3}{2}-\mathring{\kappa}\mathring{l})\Bigr{)}=\sum_{n=0}^{\infty}\sum_{\mathring{l}^{\prime}=1}^{\infty}2\mathring{l}^{\prime}\>\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}^{\prime}+\tfrac{1}{2}-\mathring{\kappa}(\mathring{l}^{\prime}-1)\Bigr{)},$
$g^{(-)}(E)$ $=\sum_{n=0}^{\infty}\sum_{\mathring{l}=1}^{\infty}2\mathring{l}\>\delta\Bigl{(}E-\hbar\omega(2n+\mathring{l}+\tfrac{3}{2}+\mathring{\kappa}(\mathring{l}+1))\Bigr{)}$ (75)
By applying the Poisson’s sum formula of the form
@@ Appendix A Berry-Tabor Formula
to the both of the double sum in Eq. (A1) and taking one of the new sums over integers using the expansion formula of trigonometric function into fractions
$\cot z=\sum_{n=-\infty}^{\infty}\frac{1}{z-n\pi}\quad\Biggl{(}\equiv\lim_{N\to\infty}\sum_{n=-N}^{N}\frac{1}{z-n\pi}\Biggr{)},\quad\mathop{\rm cosec}\nolimits^{2}z=\sum_{n=-\infty}^{\infty}\frac{1}{(z-n\pi)^{2}},$ (77)
the Berry-Tabor formula is obtaines as
$g(E)=g^{(+)}(E)+g^{(-)}(E)=\sum_{\sigma=\pm}\bar{g}^{(\sigma)}(E)+\sum_{\sigma=\pm}\sum_{L=1}^{\infty}\delta g_{L}^{(\sigma)}(E)+\sum_{N=1}^{\infty}\delta g_{N}^{(0)}(E),$ (78)
$\bar{g}^{(\pm)}(E)=\frac{\epsilon_{\pm}^{2}}{2\hbar\omega(1\mp\mathring{\kappa})^{2}},\quad\epsilon_{\pm}=\frac{E}{\hbar\omega}-1\pm\frac{1}{2}-\mathring{\kappa}$ (79)
$\delta g_{L}^{(\pm)}(E)=\frac{2}{\hbar\omega(1\mp\mathring{\kappa})^{2}}\left(\epsilon_{\pm}\cot\frac{2\pi L}{1\mp\mathring{\kappa}}\sin\frac{2\pi L}{1\mp\mathring{\kappa}}\epsilon_{\pm}+\mathop{\rm cosec}\nolimits^{2}\frac{2\pi L}{1\mp\mathring{\kappa}}\cos\frac{2\pi L}{1\mp\mathring{\kappa}}\epsilon_{\pm}\right),$ (80)
$\delta g_{N}^{(0)}(E)=-\frac{1}{\hbar\omega}\mathop{\rm cosec}\nolimits^{2}(\mathring{\kappa}\pi N)\cos(2\pi N\epsilon_{\pm}),$ (81)
which is equivalent to that in [36]. It appears to provide the trace formula (1) with contributions of POs with primitive periods $T_{\pm}=2\pi/\omega(1\mp\mathring{\kappa})$ and $T_{0}=2\pi/\omega$, but the irrational “Maslov indices” and the amplitude factors have yet to be explained by the classical dynamics.
Appendix B Other minor orbits with spin precession
@@
$U_{0}=0,\quad|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}}{M\omega},\quad|Z_{+}|^{2}+|Z_{-}|^{2}=\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}.$ (82)
@@
This condition is satisfied only by specific values of $\kappa$ and certain spin orientations for which $\kappa s_{z}$ and $\kappa s_{\perp}$ are both commensurable with $\omega$, but the orbit can exist for continuous values of energy with coefficients
$|Z_{\pm}|^{2}=\frac{1}{2}\left(\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}\pm\frac{s_{\perp}}{M\omega}\right),\quad E\geq E_{\rm bif}=\omega s_{\perp}+\kappa s_{\perp}^{2}.$
@@
$s_{z}=s_{\perp}=\frac{s}{\sqrt{2}},\quad\kappa=\frac{\omega}{24s_{z}}\fallingdotseq 0.118\,\frac{\omega}{\hbar},\quad E_{\rm bif}\fallingdotseq 0.368\,\hbar\omega.$

### v2

@@ IV.3 Bridge between circular orbits
The periodicity condition is now considered. There are three independent angular frequencies; $\omega_{p}$ for the spin precession, and $\omega\pm\kappa l$ for the orbital motion in the $z$ direction. The frequencies for the motion in $x$ and $y$ directions ($\omega\pm\kappa l\pm\omega_{p}$) are given by the combinations of the above three. Let us first consider the case where one of the coefficients $Z_{\pm}$ vanishes. In such a case, only one of the two frequencies $\omega\pm\kappa l$ should be considered in the periodicity condition. Taking into account the relation (57), let us put
$Z_{-}=0,\quad Z_{+}=\frac{s_{\perp}}{\sqrt{M\omega l}}e^{i\gamma},$ (58)
@@
$\omega+\kappa l=\frac{n}{m}\omega_{p},$ (59)
@@
Figure 4: Plot of the energy $E$ (in units of $\hbar\omega$) vs. period $T$ (in units of $2\pi/\omega$) of the POs. The vertical lines represent the diametric orbits (dotted line) and two types of circular orbits $C_{\pm}$ (dashed lines). Open circles indicate the bifurcation points (27), and the curves connected them represent the bridge orbits $(n,m)$.
@@
Each bridge orbit forms a three-parameter family since the orientation of the total angular momentum ${\bm{j}}$ and the phase $\gamma$ in Eq. (58) are arbitrary. Figure 6 displays some examples of bridge orbits $(n,m)$ for several values of $s_{z}$. As the spin vector tilts, the orbital ripples develop and eventually subside toward the spin flip.
Figure 6: Some bridge orbits $(n,m)$ with $n-m=1$, calculated for several values of $s_{z}$ (in units of $\hbar$) between $s$ and $-s$. The parameters $\kappa=0.1\,\omega/\hbar$ and $s=\hbar/2$ are used. Note that the vertical magnification is set approximately ten times the horizontal one.
@@ IV.4 Secondary bifurcation
The period of each bridge orbit $(n,m)$ always coincides with a multiple of $2\pi/\omega$ somewhere on the way from $C_{+}$ to $C_{-}$ with increasing energy. Since $\omega+\kappa l$ becomes commensurable with $\omega$ at that point, it is also commensurable with $\omega-\kappa l$ and the combination of the modes $\mu=\omega\pm\kappa l$
@@
The orbit bifurcated from the bridge $(2,1)$ is displayed in Fig. 7 for several values of energy.
@@
$E_{\rm bif}^{(n,m,k)}=\frac{n-k}{k}\frac{\omega^{2}}{\kappa}+\frac{1}{2}\kappa s^{2},$ (69)
and the orbit exists for $E\geq E_{\rm bif}$. Each orbit $(n,m,k)$ forms a four-parameter family since the orientation of ${\bm{j}}$ and the arguments $\gamma_{\pm}$ of the complex coefficients $Z_{\pm}=|Z_{\pm}|e^{i\gamma_{\pm}}$ are arbitrary. As displayed in the $E$-$T$ plots in Fig. 8, POs with the same period $T=2\pi k/\omega$ but having different geometries $(n,m,k)$ successively emerge with increasing energy.
@@
V Summary and discussion
Classical motion was investigated in the extended phase space for the system of a spherical harmonic oscillator with simple spin-orbit coupling. General solutions to the EOM and all the POs are obtained analytically, and the bifurcation scenarios are fully clarified in both generic and specific values of the spin-orbit parameter. For arbitrary value of the spin-orbit parameter, there are two kinds of degenerate two-parameter families of POs, diametric and circular ones, with frozen spin. There are two circular families C± in which spin is parallel or antiparallel to the orbital angular momentum. These circular orbits encounter successive bifurcations with increasing energy and generate three-parameter bridge PO families. On its way from one circular PO to another, each bridge encounters another bifurcation and the four-parameter PO family emerges, which has a constant period and survives for any higher energy. For specific values of spin-orbit parameter, different types of periodic solutions are also possible, for both frozen and moving spins.
Classical-quantum correspondence for this system was considered in Ref. [20]. In the direction from quantum to classical, the authors constructed the exact trace formula analogous to the Berry-Tabor formula for multiply periodic systems[26]. Quantum energy eigenvalues are given explicitly in terms of the quantum numbers as
@@ V Summary and discussion
where $n,\mathring{l},\mathring{\jmath}$ are radial, orbital angular momentum, and total angular momentum quantum numbers, respectively. Then, the level density is expressed as
$g(E)=\sum_{n,\mathring{l},\mathring{\jmath}=\mathring{l}\pm 1/2}(2\mathring{\jmath}+1)\delta(E-E_{n\mathring{l}\mathring{\jmath}}).$ (73)
By applying the Poisson’s summation formula to the above double sum, the level density is transformed into the form of summation over PO contributions as
$g(E)=\bar{g}(E)+\sum_{\rm po}A_{\rm po}\cos\left(\tfrac{1}{\hbar}T_{\rm po}E-\tfrac{\pi}{2}\mu_{\rm po}\right).$ (74)
The obtained formula includes contributions of POs with three primitive periods,
$T_{0}=\frac{2\pi}{\omega},\quad T_{\pm}=\frac{2\pi}{\omega\mp\kappa s}.$
In the direction from classical to quantum, application of the WKB method was considered in [20]. By diagonalizing the Hamiltonians with respect to the spin channels, two adiabatic Hamiltonians $H_{\pm}=H_{0}\pm\kappa s|\bm{l}|$ are obtained. Circular orbits with periods $T_{\pm}$ are found in the Hamiltonian $H_{\pm}$, respectively, which are equivalent to the circular family with frozen spin discussed in Sec. III. Straight-line orbits with period $T_{0}$ are also found in both $H_{\pm}$ but they have $\bm{l}=0$ and lie on the mode conversion point $H_{+}=H_{-}$, hence the contribution to the level density cannot be evaluated. Apart from the mode conversion problem, they noticed another difficulty in classical-quantum correspondence. Degeneracy of the PO family is related to the semiclassical order of its contribution to the level density. In general, the amplitude factor $A_{\rm po}$ in the trace formula is proportional to $\hbar^{-K/2}$ for the $K$-parameter families. In the exact trace formula, the lowest order term in the amplitude $A_{\rm po}$ for PO with periods $T_{\pm}$ is proportional to 2superscriptPlanck-constant-over-$\hbar^{-2}$ suggesting $K=4$ families. However, the circular orbits in $H_{\pm}$ form only $K=2$ families and the semiclassical order of their contributions are inconsistent.
Then, how about the Hamiltonian in the extended phase space taking the spin motion into account. The authors of Ref. [20] seem to have noticed about the existence of frozen-spin circular and diametric POs with period $T_{\pm}$ and $T_{0}$, respectively, but they didn’t discuss the matter further, probably because they do not have enough information on the classical POs. Now we have information on all the classical POs in the extended phase space. Unfortunately, the problem in semiclassical order of the contributing POs are not simply solved with the present result. The orbits having the periods $T_{\pm}$ are only the frozen-spin circular orbits that form $K=2$ families as well as in the WKB method. However, they encounter successive bifurcations with varying energy, which might enhance the contribution to the level density. Bridge orbits between circular POs have $K=3$ and their periods vary from $T_{+}$ and $T_{-}$ with increasing energy, but their contributions are not existing in the current form of the exact trace formula (74). There are two kinds of POs having the period $T_{0}$: $K=2$ family of diametric orbit with frozen spin, and $K=4$ family of ‘4D’ orbits which emerge from the second bifurcations of the bridge POs. In the exact formula, the amplitude of PO with period $T_{0}$ suggests a $K=2$ family, hence a kind of cancellation between the $K=4$ families seems to be necessary.
It is an interesting open question if the above difficulties can be overcome and a more appropriate quantum-classical correspondence can be obtained by the use of trace formula in the extended phase space with suitable uniform approximations[27, 28], and also by another way of resummation of quantum numbers from the quantum spectrum (72) to the formula (74).
Appendix Other minor orbits with spin precession
@@
$U_{0}=0,\quad|Z_{+}|^{2}-|Z_{-}|^{2}=\frac{s_{\perp}}{M\omega},$
$|Z_{+}|^{2}+|Z_{-}|^{2}=\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}.$ (75)
@@
This condition is satisfied only by specific values of $\kappa$ and certain spin orientations for which $\kappa s_{z}$ and $\kappa s_{\perp}$ are both commensurable with $\omega$, but once the orbit emerges, it can exist for any higher energy $E$ with the coefficients
$|Z_{\pm}|^{2}=\frac{1}{2}\left(\frac{E-\kappa s_{\perp}^{2}}{M\omega^{2}}\pm\frac{s_{\perp}}{M\omega}\right),$
$E\geq E_{\rm bif}=\omega s_{\perp}+\kappa s_{\perp}^{2}.$
@@
$s_{z}=s_{\perp}=\frac{s}{\sqrt{2}},$
$\kappa=\frac{\omega}{24s_{z}}\fallingdotseq 0.118\,\frac{\omega}{\hbar},\quad E_{\rm bif}\fallingdotseq 0.368\,\hbar\omega.$